How many geams of gas is produced when 65.0 mL of a 1.5M magnesium bicarbonate reacts with 35.5g of periodic acid?
Answer #1

Ans :

The balanced equation is given as :

Mg(HCO3)2 + 2HIO4 = Mg(IO4)2 + 2H2O + 2CO2

Number of mol of magnesium bicarbonate = molarity x volume (L)

= 1.5 M x 0.065 (L)

= 0.0975 mol

number of mol of acid required = 2 x 0.0975 mol = 0.195 mol

mol of acid available = mass / molar mass

= 35.5 g / 191.91 g/mol = 0.185 mol

So HIO4 is limiting reagent

Then the number of moles of CO2 gas formed will be 0.185 mol

Mass of gas = mol x molar mass

= 0.185 mol x 44.01 g/mol

= 8.14 grams